OpenStudy (anonymous):

Write an expression for the function, f(x), if it is known that f '(x) = cos(x2) and f(4) = 5. Give your answer using the form below.

OpenStudy (anonymous):

\[f(x)=\int\limits_{a}^{x}A(t)dt\]

OpenStudy (bahrom7893):

ok this is just integration. What's the integral of \[Cos(x^2)\]?

OpenStudy (anonymous):

Oh, sorry, it's f(x) = C plus the integral.

OpenStudy (anonymous):

sin(x^2)*(1/2)

OpenStudy (anonymous):

I know C=5 and A(t)=cos(t^2) But I'm having trouble finding little a to put on the bounds on the integral.

OpenStudy (bahrom7893):

Oh you need to put the bounds first?

OpenStudy (anonymous):

Yeah, you have to determine what C, a, and A(t) are. So I got that C was 5 and that A(t)=cos(t^2) due to the second fundamental theorem. But I'm stuck with finding the lower bound.

OpenStudy (anonymous):

\[f(x)=5+\int\limits_{a}^{x}\cos(t^{2})dt\]

OpenStudy (bahrom7893):

hang on let me pull up my notes on this. It's been a while.

OpenStudy (anonymous):

Okay.

OpenStudy (bahrom7893):

Wait why do you need the limits again?

OpenStudy (bahrom7893):

You don't need to know the limits, just integrate by parts: u = Cos(x^2) dv = dx

OpenStudy (anonymous):

Okay. Thank you!